OpenStudy (anonymous):

What is the set of zeros in the function? f(x)= (x^2 + 2500) x^2 - 1600)

OpenStudy (raden):

f(x)= (x^2 + 2500) (x^2 - 1600) ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

if so, the roots for f(x) is when x^2 - 1600 = 0 by using difference square formula, it can be factored be x^2 - 1600 = 0 x^2 - (40)^2 = 0 (x + 40)(x - 40) = 0 then setting each factor to zero, and solve for x

OpenStudy (anonymous):

Thanks!

OpenStudy (raden):

wait, is the roots just for real solution, or can the imaginary root be allowed ?